# Effective Learning with Node Perturbation in Multi-Layer Neural Networks

### v4

@@ 1 Introduction
Backpropagation (BP) is the workhorse of modern artificial intelligence. It provides an efficient way of performing multi-layer credit assignment, given a differentiable neural network architecture and loss function (Linnainmaa, 1970). Despite BP’s successes, it requires an auto-differentiation framework for the backward assignment of credit, introducing a distinction between a forward, or inference phase, and a backward, or learning phase, increasing algorithmic complexity and impeding implementation in (neuromorphic) hardware (Kaspar et al., 2021; Zenke and Neftci, 2021). Furthermore, BP has long been criticized for its lack of biological detail and plausibility (Grossberg, 1987; Crick, 1989; Lillicrap et al., 2016), with significant concerns again being the two separate forward and backward phases, but also its reliance on gradient propagation and the non-locality of the information required for credit assignment.
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. One such algorithm is node perturbation (NP) (Dembo and Kailath, 1990; Cauwenberghs, 1992). In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimised on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
Another proposed approach, referred to as weight perturbation (WP), is to inject noise directly into the weights of a network, rather than the activations (Werfel et al., 2003; Fiete and Seung, 2006). This method shows similarities to evolutionary algorithms and can outperform NP in special cases (Züge et al., 2021). A downside to WP is that there are many more weights than nodes in neural networks, making the exploration space larger thus slowing down learning. In fact, both NP and WP lag far behind BP in terms of efficiency. This is to be expected, however, as noise perturbations conduct a random search through parameter space, rather than being gradient directed.
@@
Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to debias layer-wise activations and thereby achieve faster learning. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects, making credit assignment more straightforward. In addition, as demonstrated by Hiratani et al. (2022), correlations in the input data lead to more bias in NP’s updates by increasing their covariance. By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivalling networks trained by BP in certain contexts.
@@ 2.1 Node perturbation and its formulations
Let us define the forward pass of a fully-connected neural network, with $L$ layers, such that the output of a given layer, $l\in{1,2,\ldots,L}$ is given by $\mathbf{x}_{l}=f\left(\mathbf{a}_{l}\right)$, where $\mathbf{a}_{l}=\mathbf{W}_{l}\mathbf{x}_{l-1}$ is the pre-activation with weight matrix $\mathbf{W}_{l}$, $f$ is the activation function and $\mathbf{x}_{l}$ is the output from layer $l$. The input to our network is therefore denoted $\mathbf{x}_{0}$, and the output $\mathbf{x}_{L}$. We consider learning rules which update the weights of such a network of the form
$\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta\Delta\mathbf{W}_{l}\,,$
@@ 2.1.1 Traditional node perturbation
$\mathbf{\tilde{x}}_{l}=f\left(\mathbf{\tilde{a}}_{l}+\bm{\epsilon}_{l}\right)=f\left(\mathbf{W}_{l}\mathbf{\tilde{x}}_{l-1}+\bm{\epsilon}_{l}\right)\,,$ (2)
where the added noise $\bm{\epsilon}_{l}\sim\mathbf{\mathcal{N}}(\mathbf{0},\sigma^{2}\mathbf{I}_{l})$ is a spherical Gaussian perturbation with no cross-correlation and $\mathbf{I}_{l}$ is an $N_{l}\times N_{l}$ identity matrix with $N_{l}$ the number of nodes in layer $l$. Note that this perturbation has a cumulative effect on the network output as each layer’s perturbed output $\mathbf{\tilde{x}}_{l}$ is propagated forward through the network, resulting in layers deeper in the network being perturbed by more than just their own added noise..
Having defined a clean and noise-perturbed network-pass, we can measure a loss differential for a NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the normalized noise vector direction and the change in loss
$\Delta\mathbf{W}_{l}^{\textrm{NP}}=\frac{\delta{\mathcal{L}}}{\sigma^{2}}\bm{\epsilon}_{l}\mathbf{x}_{l-1}^{\top}\,.$ (3)
@@ 2.1.2 Iterative node perturbation
where
$\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. Here, $\sigma^{2}$ is a scalar, equivalent to the noise variance.
@@
$\Delta\mathbf{W}_{l}^{\textrm{INP}}=N_{l}\>\delta\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||^{2}}\>\mathbf{x}_{l-1}^{\top}\,,$ (5)
@@ 2.1.3 Activity-based node perturbation
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}\,,$ (6)
where $N=\sum_{l=0}^{L}N_{l}$ is the number of units in the network, $\delta{\mathbf{a}_{l}}=\mathbf{\tilde{a}}_{l}-\mathbf{{a}}_{l}$ is the activity difference between a noisy and a clean pass in layer $l$ and $\delta{\mathbf{a}}=(\delta{\mathbf{a}_{1}},\ldots,\delta{\mathbf{a}_{L}})$ is the concatenation of activity differences.
The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix C provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of noise injected at each layer, we instead measure the total change in activation between the clean and noisy networks. This aligns our measure more closely with the directional derivative: we truly measure how the network changed in its response rather than ignoring the impact of all previous layers and their noise upon the current layer.
@@
Note that the ANP rule is now a hybridization of the directional derivative and traditional node perturbation. In the limit where the nodes are not interdependent (i.e. when we have a single layer), it is equivalent to INP and upon averaging across many samples can arbitrarily well approximate the true gradient. When used in a layered network structure, it ‘assumes’ that all nodes are mutually independent even though they are not. This is the only remaining approximation for this approach and it is desirable to allow whole network learning without direct access to the noise vectors.
@@ 2.2 Increasing NP efficiency through decorrelation
$C^{\mathrm{np}}_{kl}\approx 2C^{\mathrm{sgd}}_{kl}+\delta_{kl}\left\langle\sum_{m=1}^{k}\|\mathbf{g}_{m}\|^{2}\mathbf{I}_{k}\otimes\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}\right\rangle_{\mathbf{x}}\,,$
@@
$\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}\,,$
where $\alpha$ is a small constant learning rate and $\mathbf{R}_{l}$ is initialised as the identity matrix. For a full derivation of this procedure see (Ahmad et al., 2023).
@@
In Appendix D, Algorithm 1, we describe decorrelation using the activity-based update procedure. Appendix E illustrates the operations carried out to compute updates for the described methods.
@@ 2.3 Experimental validation
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 and CIFAR-100 datasets (Krizhevsky, 2009), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using multiple random seeds, after which performance statistics were averaged. Further experimental details can be found in Appendix F.
@@ 3.1 INP and ANP align better with the true gradient
Figure 1: Left, the angles of weight updates (relative to that computed by BP) of the NP, ANP and INP algorithms across the layers of a randomly initialized three hidden layer feedforward network. These angles are measured based on a single mini-batch of size 1000, with 10 random repeats carried out. Bar plots report average angles and error bars report standard error of the mean. Right, the number of forward passes required to compute a single update for each method. For NP and ANP two passes are required for any network update, one clean and one noisy. For INP the number of passes required is the number of layers plus one since each layer is individually perturbed and an additional clean pass is required.
Figure 1, left, shows how weight updates of the different node perturbation methods compare to those from BP in a three-hidden-layer, fully-connected, feedforward networks for the CIFAR-10 classification task. When measuring the angles between the update vectors of various methods, we can observe that the INP method is by far the most well-aligned in its updates with respect to backpropagation, followed by ANP and closely thereafter by NP. These results align with the theory laid out in the methods section of this work.
Note that for all of these algorithms, alignment with BP updates improve when updates are averaged over more samples of noise. Therefore, it is the ranking of the angles between the NP algorithms that is of interest here, not the absolute value of the angle itself, as all angles would improve with more noise samples. Appendix G shows that our results do not depend strongly on the noise variance $\sigma^{2}$, producing similar results when varying $\sigma^{2}$ by a few orders of magnitude.
Figure 1, right, shows that the node perturbation methods also differ in how many forward passes are required. Specifically, the INP method carries out an individual noise pass for each layer to isolate the impact of noise and to better approach the BP update. In the following we compare INP to the other node perturbation variants without accounting for this excess number of forward-passes, but a reader should bear in mind this drawback and consider the added computational complexity of INP.
@@ 3.2 Decorrelation improves convergence
Figure 2: Performance of node perturbation and backpropagation with and without decorrelation on CIFAR-10 when training fully-connected single-layer architecture. Curves report mean train and test accuracy. Error bars indicate the maximal and minimal accuracy obtained for five random seeds. Note that all NP methods have an equivalent formulation in a single-layer network.
To assess the impact of decorrelation on NP’s performance, we studied a single-layer network trained with NP, DNP, BP and DBP. Note that the different formulations of NP are identical in a single-layer networks where there is no impact on activities in the layer from past layers.
It appears that part of the benefit of decorrelation is due to a lack of correlation in the input unit features and a corresponding ease in credit assignment without confound. An additional benefit from decorrelation, that is specific to NP-style updates, is explained by the way in which decorrelation reduces the covariance of NP weight updates, as described in the methods section above.
@@ 3.3 Multi-layer networks can be trained effectively with node perturbation
Figure 3: Performance comparisons. Curves report mean train and test accuracy. Error bars indicate the maximal and minimal accuracy obtained for five random seeds. Top: Performance of different NP and BP formulations on CIFAR-10 when training fully-connected three-hidden-layer architectures. Middle: Performance of DANP and DINP for 3-, 6- and 9-hidden-layer networks. Bottom: Performance of different node perturbation and backpropagation variants when training a convolutional neural network on CIFAR-100.
We proceed by determining the performance of the different algorithms in multi-layer neural networks. See Appendix H for peak accuracies.
Figure  3 (Top panel) shows the various network performance measures when training a multi-layer fully connected network. Though the measured angles during training differ between these algorithms, we can see that this yields negligible differences in accuracy between the vanilla ANP and NP methods. INP does show a performance benefit relative to the other two formulations. Also, all versions of DNP are competitive with BP in terms of train accuracy, with DINP even outperforming BP. For test accuracy, BP remains the best algorithm and DBP performs by far the best on the train set. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks explored in Figure 2, meaning that DNP does facilitate multi-layer credit assignment much more than regular NP. To demonstrate that our results generalize to other datasets Appendix I reports additional results on the SARCOS dataset (Vijayakumar and Schaal, 2000).
Figure 3 (Middle panel) compares the performance of DANP and DINP for 3, 6 and 9 hidden layers. Performance for both algorithms is robust across network depth, with DINP outperforming DANP at every network depth. Though increasing the number of hidden layers beyond three did not boost performance, it also did not make the algorithms unstable, showing potential for the use of DNP in deeper networks. This is an area for future research.
To further investigate the scaling and generalization capacities of our approach, we also trained and tested convolutional architectures for image classification. Figure 3 (Bottom panel) shows that, when training a convolutional neural network on CIFAR-100, DNP, DANP and DINP outperform BP in terms of train accuracy, though BP reaches a higher test accuracy, indicating that decorrelation encourages overfitting. All decorrelated NP variants massively outperform their non-decorrelated counterparts in both train and test accuracy. Interestingly, (D)ANP outperforms (D)NP in this experiment as well.
Note that performance of all algorithms is quite low compared to CIFAR-100 benchmarks as we are using a relatively shallow convolutional neural network in combination with MSE loss, which is not the standard for classification problems with a large output space. The purpose of this experiment was not to attain competitive classification performance per-se, but a comparative study of BP and DNP under a simple training regime. These results also illustrate that the application of such perturbation-based methods can extend beyond fully-connected architectures.

### v5

@@ 1 Introduction
Backpropagation (BP) is the workhorse of modern artificial intelligence. It provides an efficient way of performing multi-layer credit assignment, given a differentiable neural network architecture and loss function (Linnainmaa, 1970). Despite BP’s successes, it requires an auto-differentiation framework for the backward assignment of credit, introducing a distinction between a forward, or inference phase, and a backward, or learning phase, increasing algorithmic complexity and impeding implementation in (neuromorphic) hardware (Kaspar et al., 2021; Zenke & Neftci, 2021). Furthermore, BP has long been criticized for its lack of biological detail and plausibility (Grossberg, 1987; Crick, 1989; Lillicrap et al., 2016), with significant concerns again being the two separate forward and backward phases, but also its reliance on gradient propagation and the non-locality of the information required for credit assignment.
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. Many focus either on greedy local learning rules or on random feedback matrices and projections, which are deemed more plausible than symmetric feedback weights (Dellaferrera & Kreiman, 2022; Frenkel et al., 2021; Kohan et al., 2018; Hinton, 2022)
The category of algorithms we focus on in this work are perturbation algorithms, such as node perturbation (NP) (Dembo & Kailath, 1990; Cauwenberghs, 1992), which require no feedback connectivity at all. In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimized on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to debias layer-wise activations and thereby achieve faster learning. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects, making credit assignment more straightforward. In addition, as demonstrated by Hiratani et al. (2022), correlations in the input data lead to more bias in NP’s updates by increasing their covariance. By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivaling networks trained by BP in certain contexts.
@@ 2.1 Node perturbation and its formulations
Let us define the forward pass of a fully-connected neural network, with $L$ layers, such that the output of a given layer, $l\in{1,2,\ldots,L}$ is given by $\mathbf{x}_{l}=f\left(\mathbf{a}_{l}\right)$, where $\mathbf{a}_{l}=\mathbf{W}_{l}\mathbf{x}_{l-1}$ is the pre-activation with weight matrix $\mathbf{W}_{l}$, $f$ is the activation function and $\mathbf{x}_{l}$ is the output from layer $l$. The input to our network is therefore denoted $\mathbf{x}_{0}$, and the output $\mathbf{x}_{L}$.
We consider learning rules which update the weights of such a network of the form
$\mathbf{W}_{l}\leftarrow\mathbf{W}_{l}-\eta\Delta\mathbf{W}_{l}$
@@ 2.1.1 Traditional node perturbation
$\mathbf{\tilde{x}}_{l}=f\left(\mathbf{\tilde{a}}_{l}+{\bf\it\epsilon}_{l}\right)=f\left(\mathbf{W}_{l}\mathbf{\tilde{x}}_{l-1}+{\bf\it\epsilon}_{l}\right)$ (2)
where the added noise ${\bf\it\epsilon}_{l}\sim\mathbf{\mathcal{N}}(\mathbf{0},\sigma^{2}\mathbf{I}_{l})$ is a spherical Gaussian perturbation with no cross-correlation and $\mathbf{I}_{l}$ is an $N_{l}\times N_{l}$ identity matrix with $N_{l}$ the number of nodes in layer $l$. Note that this perturbation has a cumulative effect on the network output as each layer’s perturbed output $\mathbf{\tilde{x}}_{l}$ is propagated forward through the network, resulting in layers deeper in the network being perturbed by more than just their own added noise.
Having defined a clean and noise-perturbed network-pass, we can measure a loss differential for a NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the loss difference and a normalized noise vector:
$\Delta\mathbf{W}_{l}^{\textrm{NP}}=\delta{\mathcal{L}}\frac{{\bf\it\epsilon}_{l}}{\sigma^{2}}\mathbf{x}_{l-1}^{\top}\,.$ (3)
@@ 2.1.2 Iterative node perturbation
where $\mathbf{v}_{k}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{k})$ if $k=l$ and $\mathbf{v}_{k}=\mathbf{0}$ otherwise. Here, $\sigma^{2}$ is a scalar, equivalent to the noise variance.
@@
$\Delta\mathbf{W}_{l}^{\textrm{INP}}=N_{l}\>\delta\mathcal{L}\frac{\mathbf{v}_{l}}{||\mathbf{v}||^{2}}\>\mathbf{x}_{l-1}^{\top}$ (5)
@@ 2.1.3 Activity-based node perturbation
$\Delta\mathbf{W}_{l}^{\textrm{ANP}}=N\>\delta{\mathcal{L}}\frac{\delta{\mathbf{a}_{l}}}{||\delta\mathbf{{a}}||^{2}}\>\mathbf{x}_{l-1}^{\top}$ (6)
where $N=\sum_{l=0}^{L}N_{l}$ is the number of units in the network, $\delta{\mathbf{a}_{l}}=\mathbf{\tilde{a}}_{l}-\mathbf{{a}}_{l}$ is the activity difference between a noisy and a clean pass in layer $l$ and $\delta{\mathbf{a}}=(\delta{\mathbf{a}_{1}},\ldots,\delta{\mathbf{a}_{L}})$ is the concatenation of activity differences. The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix C provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of noise injected at each layer, we instead measure the total change in activation between the clean and noisy networks. However, relative to BP our derived ANP rule has a biased estimation in gradient measurement. Appendix D describes this bias and shows that, when considering finite noise injection, this bias is measurable in closed form and is interpretable.
@@
Note that the ANP rule is now a hybridization of the directional derivative and traditional node perturbation. In the limit where the nodes are not interdependent (i.e. when we have a single layer), it is equivalent to INP and upon averaging across many samples can approximate the true gradient with a distinct bias based upon correlations in propagated noise. When used in a layered network structure, it ‘assumes’ that all nodes are mutually independent even though they are not. This is the only remaining approximation for this approach and it is desirable to allow whole network learning without direct access to the noise vectors.
@@ 2.2 Increasing NP efficiency through decorrelation
$C^{\mathrm{np}}_{kl}\approx 2C^{\mathrm{sgd}}_{kl}+\delta_{kl}\left\langle\sum_{m=1}^{k}\|\mathbf{g}_{m}\|^{2}\mathbf{I}_{k}\otimes\mathbf{x}_{k-1}\mathbf{x}_{k-1}^{T}\right\rangle_{\mathbf{x}}$
@@
$\mathbf{R}_{l}\leftarrow\mathbf{R}_{l}-\alpha\left(\mathbf{x}^{\star}_{l}\left(\mathbf{x}^{\star}_{l}\right)^{\top}-\textrm{diag}\left(\left(\mathbf{x}^{\star}_{l}\right)^{2}\right)\right)\mathbf{R}_{l}$
where $\alpha$ is a small constant learning rate and $\mathbf{R}_{l}$ is initialized as the identity matrix. For a full derivation of this procedure see (Ahmad et al., 2023).
@@
Figure 1 illustrates the difference between different learning rules. In Appendix E, Algorithm 1, we describe decorrelation using the decorrelated activity-based update procedure.
@@ 2.3 Experimental validation
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet datasets  (Le & Yang, 2015), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using three random seeds, after which performance statistics were aggregated. Further experimental details can be found in Appendix F.
@@ 3.1 INP and ANP align better with the true gradient
Figure 2: Left: Angles between BP’s weight update and an update calculated by NP, ANP and INP as a function of the number of noise iterations (left) and the cost of each algorithm in forward passes (right). Note that the network is not updated across the ‘Noise iterations’ dimension, but instead multiple noise samples are propagated through a fixed network and their updates averaged. This is computed for a fully-connected, 3 hidden layer, network with leaky ReLU nonlinearities. Right: Number of forward passes for NP, ANP and INP.
Figure 2, left, shows how weight updates of the different node perturbation methods compare to those from BP in a three-hidden-layer, fully-connected, feedforward network trained on synthetic data. Specifically, it is shown how each of these methods compare in the angle of their update relative to BP when you average their update over a greater and greater number of noise samples. Note that the update for Layer 1 is identical for NP and ANP.
When measuring the angles between the update vectors of various methods, we can observe that the INP method is by far the most well-aligned in its updates with respect to backpropagation, followed by ANP and closely thereafter by NP. These results align with the theory laid out in the methods section of this work. Appendix G shows the same data as a function of the number of noisy forward passes.
Note that for all of these algorithms, alignment with BP updates improves when updates are averaged over more samples of noise. Therefore, it is the ranking of the angles between the NP algorithms that is of interest here, not the absolute value of the angle itself, as all angles would improve with more noise samples. Appendix H shows that our results do not depend strongly on the noise variance $\sigma^{2}$, producing similar results when varying $\sigma^{2}$ by a few orders of magnitude.
Figure 2, right, shows that the node perturbation methods also differ in how many forward passes are required. Specifically, although all algorithms use the same dimensionality of noise the INP method requires an individual forward pass for each layer to isolate the impact of this noise and to better approach the BP update. In the following we compare INP to the other node perturbation variants without accounting for this excess number of forward-passes, but a reader should bear in mind this drawback and consider the added computational complexity of INP.
@@ 3.2 Decorrelation improves convergence
Figure 3: Top panel: Performance of (D)NP and (D)BP on CIFAR-10 when training a single-layer architecture. Bottom panel: Performance of (D)NP, (D)ANP, (D)INP and (D)BP on CIFAR-10 in a three-hidden layer network. For both figures, curves report mean train and test accuracy. Shaded areas indicate the maximal and minimal accuracy obtained for three random seeds. Note that all NP methods have an equivalent formulation in a single-layer network.
To assess the impact of decorrelation on NP’s performance, we studied a single-layer network trained with NP, DNP, BP and DBP. Note that the different formulations of NP are identical in a single-layer networks where there is no impact on activities in the layer from previous layers. Figure 3 (Top panel) indicates that, in a single-layer context, DNP outperforms NP significantly and performs slightly better than BP, with DBP performing better still.
It appears that part of the benefit of decorrelation is due to a lack of correlation in the input unit features and a corresponding ease in credit assignment without confound.
An additional benefit from decorrelation, that is specific to NP-style updates, is explained by the way in which decorrelation reduces the covariance of NP weight updates, as described in the methods section above.
@@ 3.3 Multi-layer networks can be trained effectively with node perturbation
We proceed by determining the performance of the different algorithms in multi-layer neural networks. See Appendix I, Table LABEL:tab:peakacc for peak accuracies.
Figure 3 (Bottom panel) shows the performance when training a multi-layer fully connected network. We see all decorrelated algorithms convincingly outperforming their non-decorrelated counterparts. We also see (D)INP outperform (D)ANP and (D)NP, almost keeping up with BP. Performance benefits of (D)ANP over (D)NP are not as clear in these results. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks, indicating that DNP does facilitate multi-layer credit assignment much more than regular NP, which actually performs worse in the deeper network. As an additional demonstration that our results generalize to other datasets, Appendix J reports results on the SARCOS dataset (Vijayakumar & Schaal, 2000).
Appendix  K explores the scaling of DNP, DANP and DINP to networks with 6 and 9 hidden layers as well as a convolutional neural network trained on CIFAR-10.